OpenStudy (anonymous):

x + 4 – 7x = 22 A. –3 B. 3 C. –5 D. 5

OpenStudy (anonymous):

x+4-7x=22 Subtract both sides by 4 -6x= 18 Divide both sides by - 6 Answer: -3

OpenStudy (azureilai):

Combine like ters x-7x=-6x So -6x+4=22 Then subtract 4 from both sides -6x=18 and divide by -6 x=-3

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

A :) (duh)